# Quantum entanglement, coherence and Dark Matter

### v1

@@ Appendix C Collision cross-section
As discussed in Sec. IV.2, after the transient regime, when the gas containing a large number of such systems has become dark, every two-atom system is in the anti-symmetric dark state given in equation (86). Let us then consider two such systems $A$ and $B$ entering into a collision. We choose a reference frame at rest with the center-of-mass for the two systems such that $\mathbf{R}_{A}=-\mathbf{R}_{B}=-\mathbf{R}_{0}$. For the 21 cm line of the hydrogen atom we can assume that $\left|\mathbf{P}_{A}\right|,\,\left|\mathbf{P}_{B}\right|\gg\hbar k_{r}$ irrespective of the radiation mode. To lighten the notation we omit the spatial states |$\Ket{\mathbf{R}_{A}}$, |$\Ket{\Delta\mathbf{r}_{A}}$, etc., and the vacuum |$\Ket{0}$ state to focus on the internal states, such that we can write for the colliding system
@@
Although such calculations need to be generalized or formalized, we find that two darkened entangled systems are likely to pass through each other, unimpeded, during a collision. They would not feel the presence of each other even though individual atoms may actually collide. This is due to the destructive interference resulting from the entanglement within each of the systems.

### v2

@@ Appendix D Collision cross-section
As discussed in Sec. 2.3.2, after the transient regime, when the gas has become dark, every system is in a dark state such as the one given in equation (86). Let us then consider two such systems $A$ and $B$ entering into a collision. We choose a reference frame at rest with the center-of-mass for the two systems such that $\mathbf{R}_{A}=-\mathbf{R}_{B}=-\mathbf{R}_{0}$. For the 21 cm line of the hydrogen atom we can assume that $\left|\mathbf{P}_{A}\right|,\,\left|\mathbf{P}_{B}\right|\gg\hbar k_{r}$ irrespective of the radiation mode. To lighten the notation we omit the spatial states |$\Ket{\mathbf{R}_{A}}$, |$\Ket{\Delta\mathbf{r}_{A}}$, etc., and the vacuum |$\Ket{0}$ state to focus on the internal states, such that we can write for the colliding system
@@
Although such calculations need to be generalized, we find that two darkened entangled systems are likely to pass through each other, unimpeded, during a collision. They would not feel the presence of each other even though individual atoms may actually collide. This is due to the destructive interference resulting from the entanglement within each of the systems.